### v2

# Laser Interactions with Gas Jets: EMP Emission and Nozzle Damage

Abstract
We present a theoretical model of electromagnetic pulse emission and nozzle damage from high-power laser interactions with gas jets. As hot electrons are accelerated beyond the laser channel, they produce a positively-charged plasma that expands into the ambient gas until it reaches the gas jet nozzle. A strong plasma potential leads to electrical breakdown that ionises the gas even faster than the expanding plasma ions. When contact is made between the plasma and the conducting nozzle surface, a discharge current propagates to ground and an electromagnetic pulse is radiated via an antenna emission mechanism. Damage to gas jet nozzles is caused by heating from plasma ions rather than Ohmic heating via a discharge current. Our model agrees with experimental data from experiments on proton acceleration in near-critical plasmas at the Vulcan and VEGA-3 lasers.
May 1, 2024
I Introduction
High-power laser pulses focused into dense gases are used for the acceleration of charged particles and the generation of hard x-ray radiation, but these interactions also produce undesirable secondary effects, such as nozzle damage and the emission of electromagnetic pulses (EMPs). Laser-driven EMPs are produced in the radio-frequency domain and couple to motors, computers and other electronic equipment. On the VEGA-3 laser system [1] at CLPU, EMPs have been responsible for valve malfunctions and gas leaks from jet nozzles; they are also known to enter diagnostics and oscilloscopes - ruining measurements of charged particle emission.
Megahertz- and gigahertz-frequency EMPs are generated in a variety of high-power laser experiments when hot electrons are expelled from the target and oscillating currents are excited in the target mount and surrounding chamber [2]. Previous research [3, 4, 5, 6, 7] has focused on laser interactions with solid targets, where the EMP amplitude is known to be the highest. The few measurements available for high-density gas jet targets [8, 9], however, suggest that EMP emission is significant. If the EMP amplitude from gas jets scales with laser energy and intensity as for solid targets, these EMP fields will increase with a new generation of ultra-intense, high-repetition-rate laser systems [10].
A second important concern of the laser-gas interaction is damage to the gas jet nozzle. Many of the most exciting applications of laser-gas research rely on high shot rates and a reproducible gas density profile to generate bright, high-fluence sources of energetic ions [11, 12, 13, 14, 15, 16], electrons[17, 18] or x-rays [19, 20]. Repeated melting of the gas nozzle therefore represents a serious and expensive hindrance to this research. On energetic systems like the Vulcan Petawatt laser [21, 22] at the Rutherford Appleton Laboratory, a single shot on a dense gas target is sufficient to destroy the gas nozzle completely (see Fig. 1). For systems operating at lower energy, lower gas density and shorter pulse duration, nozzle damage is more progressive but still leads to significant smoothing of the gas density profile [23, 24, 25], degradation of the laser interaction and reduced data reproducibility. Understanding the nozzle damage mechanism is therefore important for future laser-gas applications.
Figure 1: Gas jet nozzle used during an experiment at Vulcan Target Area Petawatt (a) before and (b) after a full-power laser shot. Melting can lead to occlusion of the nozzle aperture or even total rupture of the material. Images reproduced from Hicks et al. [26] with permission.
To date, no theoretical framework has been proposed for the emission of EMP from gas jets and this makes it difficult to estimate the severity of its impact on new laser systems. In Sec. II and Sec. III.1-III.3, we present a model of laser-plasma expansion in a gas, where ionisation is mainly caused by ions streaming from the laser focus (anode) to the nozzle (cathode). In Sec. III.4, however, we show that electrical breakdown induced by a strong plasma-nozzle potential can fully ionise the gas on shorter timescales of 10-20 ps. EMP is then emitted similarly to solid-target interactions, where a discharge current propagates to ground (the chamber) along a cm-scale antenna (the nozzle or target support), emitting radiation at gigahertz frequencies. According to our model, the EMP amplitude for Vulcan Petawatt and VEGA-style interactions can reach $\sim 100$ $\mu$T or $\sim 100$ kV/m at a distance $\sim 1$ m from the target. The theoretical model is supported in Sec. VI by the results from two experiments on the VEGA-3 laser, which show reasonable agreement with the discharge time, magnetic field spatial profile and accumulated target charge.
In Sec. V, we consider how the nozzle may be damaged using the plasma expansion model. Two mechanisms are considered: (i) a kA-level discharge current and (ii) ion collisional heating. We show that damage to gas jet nozzles is more likely caused by the impact of plasma ions than Ohmic heating of the nozzle surface. Finally, in Sec. VII, we present ideas for how the EMP emission model can be reliably benchmarked with simulations and dedicated experiments. The practical impact of this work is broad: allowing scientists to reduce damage to expensive gas jet nozzles and minimize the electrical disruption of equipment. It likewise represents a rich seam of more fundamental research, connecting the physics of laser-target charging, ionisation, ion acceleration, high-voltage breakdown and antenna emission.
II Expansion of a Laser-Plasma in a Gas Jet
We consider an expanding plasma created as a laser pulse propagates through a high-density gas. The laser pulse ionises the gas and creates a plasma channel. Plasma electrons are heated by the laser radiation to relativistic energies, and some of them escape the channel and leave it positively charged. This charged plasma cylinder then expands into the surrounding gas or plasma, depending on the efficiency of the ionisation mechanisms discussed in Sec. III. If the plasma expands into gas, the gas is ionised until contact is made with the conducting tip of the gas nozzle. Once contact is made and the plasma is connected directly to ground, a discharge is triggered and EMP radiation is emitted. Nozzle damage is determined variously by the amount of energy stored in the plasma, the strength of the discharge and the nozzle material.
Starting from some characteristic gas and laser parameters, we first explore the physics of plasma expansion, electrical discharge and EMP emission. By way of example, we consider the Vulcan-Petawatt laser, with pulse energy $E_{L}=200$ J, duration $\tau_{L}=600$ fs and wavelength $\lambda_{L}=1.053\,\mu$m [26]. The laser beam is focused using $f_{\#}=3$ optics to a spot with full width at half maximum (FWHM) $w_{L}=3.9\,\mu$m. Only 40% of the total laser energy is contained within the FWHM, for a maximum on-target intensity of $I_{L}\simeq 2.9\times 10^{20}$ W cm-2. The spatial extent of the laser channel in the gas depends on the wavelength of the laser and the gas density. For typical gas pressures $p_{\rm gas}$ ranging from 0.03 to 30 bar, the atomic/molecular densities in the laser focal region are in the range $n_{\rm gas}\sim 10^{18}-10^{21}$ cm-3. The maximum electron density in a fully ionised plasma is smaller than the critical density, so the laser can propagate through the gas.
The Rayleigh length of a diffraction-limited Gaussian beam in our example is $Z_{R}=\pi w_{L}^{2}/\lambda_{L}\simeq 44\,\mu$m. Since the gas jet width is usually larger than $Z_{R}$ and the beam is susceptible to relativistic self-guiding, the laser produces a channel a few times longer than $Z_{R}$ that can be viewed as a $l_{p}\simeq 0.5$ mm-long plasma cylinder created at a typical height of $h\simeq 0.5$ mm above the nozzle. We take the initial radius of the plasma channel to be a factor of a few times larger than the laser focal radius, $r_{p}=20\,\mu$m, to reflect the average size of the channel as the laser deposits energy over the full length of its path through the gas. We assume that laser absorption in the gas jet is $\eta_{L}\sim 90$%, with 40% transferred to hot electrons. This is consistent both with Particle-in-Cell (PIC) simulations[26, 27] and experimental interferograms [28], which show that a high-power laser pulse deposits a large fraction of its energy in the gas jet and leaves behind it a plasma channel several tens of microns wide.
Parameter Vulcan-TAP VEGA-3
Gas H2 He
$\lambda_{L}$ [$\mu$m] 1.05 0.8
$E_{L}$ [J] 80 8
$\tau_{L}$ [fs] 600 75
$f_{\#}$ 3 10
$w_{L}$ [$\mu$m] 3.9 9.8
$I_{L}$ [W cm-2] $2.9\times 10^{20}$ $3.6\times 10^{19}$
$a_{0}$ 15 4.1
$\eta_{L}$ 0.9 0.9
$\eta_{\rm las\to h}$ 0.4 0.1
$p_{\rm gas}$ [bar] 15 4.5
$h$ [$\mu$m] 500 500
$r_{p}$ [$\mu$m] 20 20
$V_{p}$ [cm3] $6\times 10^{-7}$ $6\times 10^{-7}$
$n_{e}$ [cm-3] $8.1\times 10^{20}$ $2.4\times 10^{20}$
$T_{e}$ [keV] 280 140
$T_{h}$ [MeV] 5.0 1.1
$\Phi_{p}$ [MV] 25 4.4
$Q$ [nC] 230 42
Table 1: Representative experimental parameters - as described in the text and Refs. [26, 24] - and estimates of plasma characteristics according to the model presented in Sec. II. The plasma radius $r_{p}$ and the laser energy fraction converted to hot electrons $\eta_{\rm las\to h}$ are estimated from dedicated PIC simulations.
The parameters of the experiments considered in this paper are given in Tab. 1 along with predictions from our model. Depending on the nozzle design and envisioned application, the laser pulse can be sent parallel to the surface of the nozzle tip at a height varying from a few tenths of a millimetre to a few millimetres. The width of the laser channel also depends on the laser focusing optics, laser power, and gas density profile along the laser axis. Figure 2 gives a schematic overview of the situation.
Figure 2: Schematic of the nozzle, gas and the cylindrical plasma channel formed by the laser. Here, $r_{p}$ is the radius of the plasma cylinder, $h$ is the separation between the channel and the nozzle, and $\Phi_{p}$ is the electric potential. The laser is directed into the plane of the page, along the axis of the plasma. Red crosses indicate a region of positive charge inside the laser-generated channel. Three curved arrows sketch the geometry of the electric field, E.
II.1 Electric Charging of the Plasma
Following the available observations and numerical simulations [29, 30, 31, 32], we assume that the plasma in the laser channel is fully ionised. The absorbed laser energy $\eta_{L}E_{L}$ is transferred to electrons with an energy distribution approximated by a two-temperature relativistic distribution:
$f_{e}(\gamma)=n_{e}f_{\rm MJ}(\gamma,\Theta_{e})+n_{h}f_{\rm MJ}(\gamma,\Theta_{h}),$ (1)
where $f_{\rm MJ}(\gamma,\Theta)$ is the Maxwell-Jüttner function[33]
$f_{e}(\gamma,\Theta)=(\gamma/\Theta)\,K_{2}(1/\Theta)\sqrt{\gamma^{2}-1}\exp(-\gamma/\Theta)$
depending on the dimensionless temperature $\Theta=T/(m_{e}c^{2})$, the modified Bessel function $K$ of the second kind and the electron relativistic factor $\gamma=1+\varepsilon/(m_{e}c^{2})$. A fraction $\eta_{\rm las\to h}$ of the absorbed laser energy is transferred to a population of hot electrons with density $n_{h}$ and temperature $T_{h}$, while the remaining energy goes to electrons in the plasma bulk, with density $n_{e}$ and temperature $T_{e}$. Here, we use subscripts “h” and “e” to refer to the hot and thermal electron populations respectively.
The bulk electron density is defined by the gas density ($n_{e}=Zn_{\rm at}$) and the bulk temperature is defined by the plasma volume and deposited energy via:
$C_{e}(T_{e})n_{e}T_{e}=(1-\eta_{\rm las\to h})\eta_{L}E_{L}/V_{p},$ (2)
where $C_{e}(T_{e})=3-[1-K_{1}(1/\Theta_{e})/K_{2}(1/\Theta_{e})]/\Theta_{e}$ is the heat capacitance of a relativistic electron gas. For the parameters of the Vulcan experiment given in Tab. 1, a plasma channel of radius $r_{p}=20\,\mu$m and length $l_{p}=0.5$ mm has volume $V_{p}=\pi r_{p}^{2}l_{p}\simeq 6\times 10^{-7}$ cm3. A hydrogen gas pressure of 15 bar corresponds to an atomic gas density $n_{\rm at}\simeq 8.1\times 10^{20}$ cm-3 at a distance $h\simeq 0.5$ mm above the nozzle, so we have $N_{e}\simeq 5.1\times 10^{14}$ electrons. Equation (2) yields an effective electron temperature $T_{e}=250$ keV for the total bulk energy of 43 J.
To estimate the hot electron density, we prescribe a hot electron temperature of $T_{h}=5.0$ MeV following the ponderomotive scaling[34], $(\sqrt{1+a_{0}^{2}}-1)m_{e}c^{2}$. The hot electron density can then be defined knowing the fraction of energy transferred to hot electrons $\eta_{\rm las\to h}$:
$C_{e}(T_{h})n_{h}T_{h}=\eta_{\rm las\to h}\eta_{L}E_{L}/V_{p}.$
Knowing the hot electron temperature and assuming that $\eta_{\rm las\to h}=0.4$ on Vulcan, this equation gives a hot electron fraction $n_{h}/n_{e}\simeq 2.5$% with total energy 29 J.
The electric potential of the plasma channel is determined by how many hot electrons can overcome the plasma potential barrier. Though electrons from the plasma bulk will also contribute to the ejected charge, their contribution is less than $10\%$ because a significant fraction of the available laser energy is transferred to hot electrons. The electric potential $\Phi_{p}$ at the surface of the plasma cylinder with respect to the grounded nozzle can be estimated as
$\Phi_{p}=Q/C,$ (3)
where $Q$ is the charge of the plasma and $C=2\pi l_{p}\epsilon_{0}/\ln(h/r_{p})$ is the capacitance. The capacitance follows from the potential of a charged cylinder, radius $r_{p}$, length $l_{p}$, situated a distance $h$ from an infinite conducting plane. Only electrons with energies $\varepsilon_{e}>e\Phi_{p}$ can escape plasma to the ground. Assuming, as mentioned above, that hot electrons have an exponential distribution in energy with a temperature $T_{h}$, the fraction of escaped hot electrons is [35]:
[…]
where $\psi=e\Phi_{p}/(m_{e}c^{2})$ is the dimensionless potential. The plasma charge is then $Q=e\delta n_{h}V_{p}$ and, substituting this expression into Eq. (3), we have the following equation for the dimensionless potential $\psi$:
$\psi=\frac{r_{p}^{2}\omega_{ph}^{2}}{2c^{2}}\,\frac{\delta n_{h}}{n_{h}}\,\ln\frac{h}{r_{p}},$ (5)
where $\omega_{ph}=(e^{2}n_{h}/(m_{e}\epsilon_{0}))^{1/2}$ is the hot electron plasma frequency. The solution of this equation for typical parameters $h/r_{p}=250$ and $r_{p}\omega_{ph}/c=18$ gives $\Phi_{p}=24.8$ MV. Knowing the plasma potential and the plasma capacitance $C\simeq 0.0086$ pF, we find the charge ejected from the plasma column, $Q=C\Phi_{p}\simeq 234$ nC. This corresponds to a fraction of escaped electrons $\delta n_{h}/n_{e}\approx 2.8\times 10^{-3}$ and is consistent with the experiment [36]. Since the time required for a hot electron to cross the channel is relatively short ($\sim 2r_{p}/c=0.13$ ps), the time of plasma charging is determined by the propagation time of the laser pulse through the gas, $\sim 3$ ps.
The total energy of escaped electrons
[…]
is split between the electrostatic energy of the charged plasma $\mathcal{E}_{\rm es}=Q^{2}/2C\simeq 3.2$ J and the kinetic energy of the escaped electrons $\simeq 4.2$ J.
II.2 Ion Acceleration in Expanding Plasma
Following a short phase of gas ionisation, electron heating and plasma charging, the plasma expands and cools down. The time required for the plasma to expand from the laser-generated channel to the nozzle defines the plasma discharge time. In Sec. IV, we show that the discharge time determines the energy and amplitude of the radiated EMP. Plasma x-ray emission and laser-accelerated fast electrons and ions, as shown in Sec. III, partially ionise the gas and provide seed electrons for a discharge via electrical breakdown.
After the end of the laser pulse, there is no more energy supply and the plasma expands adiabatically under the electron thermal pressure. For the plasma parameters given above, the pressure $p_{e}=n_{e}T_{e}\simeq 330$ Mbar is much higher than the ambient gas pressure, so the plasma expands freely with electrons transferring their energy to ions. The plasma charge is conserved during the expansion phase and the potential decreases logarithmically as the plasma radius increases (see Eq. (5)). The plasma expansion can be described in some special cases by a self-similar rarefaction wave model [37, 38]. More detailed analysis is performed for a spherical plasma expansion in Ref. [39]. The ion energy distribution depends on the density profile and on the ratio of the plasma radius to the Debye length. For our purposes, it is sufficient to estimate the average ion kinetic energy by dividing the deposited laser energy by the number of ions in the plasma. For a hydrogen plasma, it is 1.5 times the bulk electron temperature $\varepsilon_{i}\simeq 0.4$ MeV. This corresponds to an ion velocity of $v_{i}\simeq 9\,\mu$m/ps, and an ion expansion time to the nozzle $t_{\rm exp}\simeq 60$ ps. The total energy carried by these ions is of order the bulk electron energy - that is, about $(1-\eta_{\rm las\to h})\eta_{L}E_{L}\sim 40$ J for the Vulcan experiment.
In addition to ions accelerated by bulk electron pressure in the expanding plasma, some ions are accelerated to energies of a few MeV in the electrostatic field created by the escaped electrons. Hicks et al. [26] report on the number of such fast protons accelerated in the radial direction as 30 nC/sr. Assuming an emission solid angle of $1-2$ sr, the total charge of ejected protons is about 60 nC and their overall energy about 0.2 J. These values are consistent with 230 nC charge of ejected electrons and a corresponding energy of $\sim 4$ J, as estimated above.
III Gas Ionisation by Laser-Plasma Products
The level of EMP emission depends on the intensity of the discharge current and, consequently, the resistance of the plasma that connects the laser channel to the nozzle. There are four processes of gas ionisation: (i) photoionisation from UV and x-rays produced by the bremsstrahlung of hot electrons in the plasma column, (ii) collisional and field ionisation from fast electrons accelerated in the laser channel, (iii) collisional ionisation from fast ions emanating from the laser channel and (iv) ionisation by electron avalanche caused by a high plasma-nozzle electric potential.
III.1 X-ray Photoionisation of the Gas
The bremsstrahlung emission of electrons in the laser-created plasma covers a broad range of photon energies up to the electron kinetic energy. Assuming the electron energy distribution is characterized by an effective temperature $T_{e}$, the power per unit volume of bremsstrahlung emission from a hydrogen plasma can be written as [40]:
$P_{\text{Brem}}=1.69\times 10^{-32}n_{e}^{2}T_{e}^{1/2}\,{\rm W\leavevmode\nobreak\ cm}^{-3},$
where the electron density is in cm-3 and the electron temperature is in eV. The radiative contribution of bulk electrons dominates that of hot electrons because $n_{e}T_{e}\gtrsim n_{h}T_{h}$ under the present conditions. For an electron density of $n_{e}=8\times 10^{20}$ cm-3 in fully ionised plasma and a temperature $T_{e}\simeq 250$ keV, we have $P_{\text{Brem}}=6\times 10^{12}$ W cm-3. Accounting for the plasma volume of $V_{p}\sim 6\times 10^{-7}$ cm3 and lifetime $t_{\rm exp}\sim 60$ ps, the total radiated energy is about $\mathcal{E}_{\rm Brem}=P_{\rm Brem}V_{p}t_{\rm exp}\simeq 0.2$ mJ. The spectrum of emission is flat for photons with energy $\varepsilon_{x}\lesssim T_{e}$ and decreases exponentially with a temperature $T_{\rm rad}\simeq T_{e}$ for higher energies.
The mean free path of photons with energies above 3 keV in hydrogen is approximately constant at $\lambda_{x}\rho_{\rm gas}\sim 3$ g cm-2[41]. For a gas jet density of $\rho_{\rm gas}\sim 1.3$ mg cm-3, the absorption length is 20 m and the probability of gas ionisation over a millimetre distance is negligible. The mean free path of lower energy photons strongly decreases as the energy decreases [41], however. It can be approximated by $\lambda_{x}\rho_{\rm gas}\sim 10^{-10}\varepsilon_{x}^{3}$ g cm-2 for hydrogen, where $\varepsilon_{x}$ is the photon energy in eV. The mean free path of photons with the energy of $15-20$ eV - comparable to the hydrogen ionisation potential $U_{\rm ion}=13.6$ eV - is therefore just a few microns.
The photons with energy about $\varepsilon_{h}\sim 100-200$ eV produce the most efficient photoionisation because their mean free path is comparable to the plasma-nozzle distance. The fraction of bremsstrahlung energy carried by these photons is $\varepsilon_{h}/T_{e}\sim 10^{-3}$ that is, about $0.2\,\mu$J. Knowing the energy required to ionise a hydrogen atom and the volume of gas between the plasma and the nozzle, $V_{\rm gas}\sim 4\times 10^{-4}$ cm3, the gas ionisation level can be estimated as $\mathcal{E}_{\rm Brem}\varepsilon_{h}/T_{e}U_{\rm ion}V_{\rm gas}\sim 1.7\times 10^{14}$ cm-3. This corresponds to an ionised fraction of approximately $3\times 10^{-7}$. The photoionisation time of $\sim 60$ ps is defined by the time of bremsstrahlung emission, which is the lifetime of the hot plasma channel.
III.2 Ionisation by Hot Electrons
Hot electrons also contribute to the ionisation process. These electrons are those escaping the plasma potential barrier and their characteristics are estimated in Sec. II. The stopping power of electrons in hydrogen gas for this range of energies of a few MeV is $\sim 4$ MeV cm2/g  [42], so a hot electron loses $\sim 260$ eV over the distance $h\simeq 0.5$ mm. The energy of secondary electrons is comparable to the ionisation potential of the gas [43], which means that one primary electron creates $\sim 20$ secondary electrons in the gas volume. Using the number of escaped electrons estimated in Sec. II.1, $N_{\rm esc}\simeq 1.4\times 10^{12}$, the density of secondary electrons produced by collisional ionisation is $\sim 7.5\times 10^{16}$ cm-3. This corresponds to an ionised fraction of $\sim 10^{-4}$, which is produced in a short time of hot electron emission of a few ps.
[…]
III.3 Ionisation by Plasma Ions
The stopping power of a representative 0.4 MeV proton in hydrogen is 1200 MeV cm2/g  [45]. The energy needed for the creation of one secondary electron by a fast ion is of order the ionisation potential, $U_{\rm ion}$, and it varies weakly with the ion energy[46, 47]. A 0.4 MeV proton passing through hydrogen gas therefore creates $\sim 6000$ secondary electrons over the 0.5 mm distance between the plasma and the nozzle, losing about 0.1 MeV of its kinetic energy. In a gas volume of $\sim 4\times 10^{-4}$ cm-3, this corresponds to complete ionisation of the ambient gas. This is, however, a relatively long process that takes $\sim 60$ ps.
The contribution of fast ions to gas ionisation is smaller by more than three orders of magnitude, but their velocity is about three times higher. An ionisation level of $\sim 3\times 10^{-3}$ is therefore produced prior to plasma expansion, during the time of flight of the fast ions, which is about 20 ps.
In conclusion, photoionisation of the gas jet is low and can be neglected. Instead, ionisation by charged particles is produced in three steps: first, at the level of $10^{-4}$ by fast electrons in a few ps, then at the level of $10^{-3}$ by fast ions within 20 ps, then finally by the expanding plasma on a timescale of 60 ps. Free electrons produced during the first two steps provide the seed for a discharge by electrical breakdown, as discussed in the next section.
III.4 Gas Ionisation by Electrical Breakdown
When hot electrons are ejected from the laser plasma, it becomes positively charged with respect to the nozzle. The electric field between the plasma and nozzle may then break down the gas and propagate a discharge. The discharge regime is determined by the product of the gas pressure and electrode separation $p_{\rm gas}h$. Under Vulcan-like conditions, with $p_{\rm gas}h\gtrsim 1$ bar cm, avalanche ionisation will occur if the potential is above the breakdown value of $20-30$ kV  [48]. This is much smaller than the plasma potential of $\sim 25$ MV estimated in Sec. II.
In the avalanche process, the electron density increases exponentially in time from the seed level $n_{e0}/n_{\rm at}\sim 10^{-4}-10^{-3}$ created by fast electrons or ions, $n_{e}(t)=n_{e0}\,\exp(\nu_{\rm ion}t)$, where $\nu_{\rm ion}$ is the characteristic ionization rate. According to Refs. [48, 49], the ionisation rate is $\nu_{\rm ion}\sim 10^{12}$ s-1 for Vulcan-like conditions and the gas can be fully ionised in a few picoseconds. However, the plasma potential decreases with time for two reasons: first, it decreases logarithmically with time because of plasma expansion according to Eq. (5) and second, because the energy deposited during the ionisation process is extracted from the electrostatic energy. The decrease in electrostatic energy is related to the ionisation loss, $\Delta\mathcal{E}\simeq U_{\rm ion}n_{e}V_{\rm gas}$, and the discharge stops when all the available electrostatic energy $\mathcal{E}_{es}$ has been exhausted. This relation defines the maximum ionisation that can be achieved by avalanche breakdown, $n_{e,\max}\sim 4\times 10^{21}$ cm-3, which is larger than the atomic gas density.
In conclusion, avalanche breakdown is initiated from the seed level of ionisation produced by fast electrons or fast ions on a $10-20$ ps timescale. It can fully ionise the gas under the conditions of the Vulcan experiment before the expanding plasma reaches the nozzle.
III.5 Plasma Resistance
The degree of gas ionisation determines the plasma resistance and the discharge current that can be supported. We estimate the plasma resistance as $R=\eta h/A$, where $\eta$ is the plasma resistivity and $A\approx 2r_{p}l_{p}$ is the cross-sectional area of the ionised gas between the nozzle and the laser channel. The plasma resistivity in a fully ionised plasma is dominated by electron-ion collisions and can be estimated from the Drude formula:
$\eta=m_{e}\nu_{ei}/e^{2}n_{e},$ (6)
where $\nu_{ei}$ is the electron-ion collision frequency. Assuming the electron temperature to be of the order of the ionisation potential, we find $\nu_{ei}\sim 10^{14}$ s-1 and the plasma resistivity is $\eta\simeq 5\times 10^{-6}\,\Omega$ m. This corresponds to a very small resistance of order $0.1\,\Omega$. A path to ground is established when the gas is fully ionised; the potential drop across the nozzle is of comparable magnitude to the plasma potential and a discharge current is produced. The accumulated plasma charge and discharge time puts a limit on the maximum discharge current of order $\sim 10$ kA, as discussed in the next section.
IV EMP emission
Here, we estimate the characteristics of EMP emission. The theoretical picture is as follows: the plasma supplies a current limited by the nozzle resistance and the available charge and radiation is emitted from an antenna made out of the conducting parts of the nozzle. Considering the nozzle as a quarter-wavelength dipole, the emission frequency depends on the nozzle length $h_{d}$ as $\nu_{\rm emp}=c/4h_{d}$. For $h_{d}=3$ cm, the emission frequency is 2.5 GHz.
The EMP energy is limited by the available electrostatic energy in the plasma, which is of order a few joules. Modelling the plasma-nozzle system as a dipole, the total emitted energy, according to the textbook by Jackson [50], depends on the charge $Q$ and the emission frequency:
$\mathcal{E}_{EMP}\approx 0.1Z_{0}Q^{2}\nu_{\rm emp}$ (7)
where $Z_{0}=377$ $\Omega$ is the vacuum impedance. For 230 nC accumulated charge and an emission frequency of 2.5 GHz, the relation yields $\sim 5$ mJ of EMP energy and an average current at the antenna frequency[51] of $J_{\text{emp}}=Q\nu_{\rm emp}\sim 580$ A. EMP emission is delayed with respect to the laser: first, because it takes time to ionise the gas and second, because the current needs time to propagate along the antenna.
Knowing the amplitude of the discharge current, we can also estimate the amplitude of the emitted magnetic field at a distance $r$ as $B\simeq\mu_{0}J_{\text{emp}}/2\pi r$. This is equivalent to the maximum B-field (measured over all possible emission angles) in the far-field of a dipole antenna[51]. For a nozzle of height $h_{d}=3$ cm, the peak magnetic field amplitude is $\sim 120$ $\mu$T at a distance of $r=1$ m. The gas discharge is thus expected to produce an electromagnetic pulse with electric field of amplitude $E=cB\approx 36$ kV/m (in the plane wave approximation). This value is higher than the electromagnetic susceptibility threshold for electronics [52].
V Nozzle Damage Mechanism
We identify two mechanisms [^1] UV, soft x-rays and thermal electrons can also couple to the nozzle and produce heating. Hard x-rays will deposit their energy volumetrically and are less important. The energy in UV and x-rays is relatively small, however (see Sec. III.1), while the thermal electrons travel with the ions and have a lower stopping power[42], producing heating in conjunction with the ions. that could lead to nozzle damage in high-power laser experiments with gas jets: Ohmic heating from a discharge current or bombardment by plasma ions. To evaluate which process is more likely to induce melting, we calculate the energy required to melt a given mass of nozzle material.
We consider copper as a representative nozzle material, with a heat capacity $C_{V}=0.39$ J/gK, a melt temperature of 1360 K and a latent heat of melting 0.86 kJ/g. To heat 1 g of Cu to the melting temperature - and to melt it - one needs a total energy of 1.3 kJ. Similar estimates apply to aluminium and iron. The total available electrostatic energy in the plasma of 3 J can melt a maximum of 2 mg Cu. By contrast, the kinetic energy of the expanding plasma is more than 10 times larger. The plasma can therefore cause more damage to the nozzle.
V.1 Energy Released through Ohmic Heating
The energy released by Ohmic heating is $J^{2}Rt_{p}$, where $J=Q/t_{p}$ is the discharge current. The resistance $R$ is calculated from the metal resistivity $\eta_{d}$ and the current-carrying surface $A_{d}$ and nozzle length $h_{d}$. Since the discharge is short, the current is confined within a skin depth, $\delta=\sqrt{\eta_{d}/(\nu_{\rm emp}\mu_{0})}$. For Cu resistivity $\eta_{d}=1.7\times 10^{-7}\,\Omega$ m and the frequency $\nu_{\rm emp}\approx 2.5$ GHz, the skin depth is $\delta\sim 10\leavevmode\nobreak\ \mu$m. Assuming the current flows along $h_{d}=3$ cm nozzle with a radius of $r_{d}=5$ mm, the current-carrying cross-sectional area is $A_{d}\sim 3\times 10^{-3}$ cm2 and the resistance is $R=\eta_{d}h_{d}/A_{d}\approx 0.01\,\Omega$. The energy released by Ohmic heating with a kA-level current over $t_{p}\sim 2h/c\approx 3$ ps is therefore $\lesssim 0.1\,\mu$J.
Will the conducting layer melt if it is supplied with $0.1\,\mu$J of heat energy? The volume of heated material is $h_{d}A_{d}\sim 10^{-2}$ cm3, which gives a heated mass of $\sim 80$ mg. The available electrostatic energy is not sufficient to heat such a large mass, so it is unlikely that the melting of gas nozzles is caused by Ohmic heating from a discharge current.
V.2 Heating caused by Plasma Deposition
Plasma ions are the primary source of nozzle heating. Assuming that $N_{i}$ ions in the plasma are expanding isotropically, the angular ion flux is $N_{i}/4\pi$. The solid angle subtended by the nozzle is given by $2\pi(1-\cos\alpha)$, where $\tan\alpha=r_{d}/h$. For $h=0.5$ mm and $r_{d}=5$ mm, the solid angle is approximately $2\pi$ and the number of ions incident on the nozzle is around half the ion population.
[…]
Based on the estimates of plasma density and volume from Sec. II.1, there are $N_{i}\sim 5\times 10^{14}$ ions at energy $\sim 0.4$ MeV for a combined energy of 40 J. Half of these ions strike the nozzle and deposit an energy of 250 kJ/g. This is more than 100 times higher than the energy required to melt Cu, so plasma heating melts the nozzle easily.
The same conclusion applies to a ceramic nozzle, even with the greater heat resistance. Consider SiC with a melting temperature of 3100 K, 370 J/g latent heat of melting and a specific heat capacity of $0.67-1.4$ J/gK for temperatures between $300-4000$ K [54]. Using these values, we find that $\sim 4$ kJ/g is required to melt SiC. According to the NIST proton stopping catalogue, the stopping power of a 0.5 MeV proton in SiO2 is 290 MeV cm2/g [45]. For a density of 3.2 g cm-3, the rate of energy deposition is 1000 MeV/cm, and the deposition depth is $\sim 5\,\mu$m, for a heated volume of approximately $2\times 10^{-5}$ cm3 and a heated mass of $0.06-0.1$ mg. This is consistent with SRIM Monte Carlo simulations [55] of 0.4 MeV protons incident on SiC, which give a $4\,\mu$m range. 20 J of incident ion energy deposited in a few microns corresponds to heating of $\sim 2$ MJ/g - more than sufficient to melt the ceramic.
This conclusion is supported by a gas jet experiment on the VEGA-3 laser [23]. The laser was focused to an intensity of $2.2\times 10^{19}$ W cm-2 in a He (97%) and H2 (3%) gas mix with peak atomic densities of order $\sim 10^{21}$ cm-3 (see Sec. VI.2 for more details of the experiment). Tungsten nozzles suffered progressive damage over tens of shots when the laser was focused at a distance $\sim 900\,\mu$m, whereas a UV fused silica nozzle was destroyed when the laser was focused at a distance of $400\,\mu$m from the nozzle surface.
We can apply this model of nozzle damage to a Vulcan-style experiment. The amount of energy deposited by the ions is sensitive to the number of ions and their average energy. Taking a vertical gas density profile from measurements by Hicks et al. [26], we estimate the number of ions generated in the laser channel when the laser is focused at different heights above the gas jet. We assume a value of $N_{i}$ at a fixed height of $160\,\mu$m, where the density is maximal, and scale it by the normalised vertical density profile as the laser focus is shifted to different heights above the nozzle. Following the method outlined earlier, we then calculate the ion flux on the nozzle surface and the amount of energy deposited when the laser is focused at different distances from the nozzle. Hicks et al. observed that metallic nozzles were instantly destroyed at a nozzle-focus distance of $400\,\mu$m and virtually undamaged when they were separated by 2 mm. Since the melt threshold for Cu is 1.4 kJ/g, this suggests that the ion energy deposition must reach 2 kJ/g when the laser is focused somewhere between these two distances. We therefore have bounds on the total number of 0.4 MeV plasma ions produced by the laser at a height of $160\,\mu$m, which satisfies $4\times 10^{12}<N_{i}<2\times 10^{14}$. The product of the initial plasma volume and peak gas density gives a maximum electron number of $\sim 5\times 10^{14}$, which may be considered an upper limit on the number of accelerated ions.
Figure 3: Ion energy deposited per mass of nozzle material (Cu or Fe) as a function of laser-nozzle separation for the number of ions $N_{i}=4\times 10^{12}$ (a) and $2\times 10^{14}$ (b) normalised to the plasma height of $160\,\mu$m. The red dot represents the theoretical melt threshold for a Cu nozzle. Red and green vertical dashed lines represent observed distances where a steel nozzle was destroyed and survived, respectively. The ion energy is 0.5 MeV.
The situation is illustrated graphically in Fig. 3. The left panel shows the ion energy deposited in the nozzle as a function of distance, assuming $4\times 10^{12}$ plasma ions at 0.5 MeV when the laser is focused at a height of $160\,\mu$m. The right panel shows the energy deposition for $2\times 10^{14}$ plasma ions at a height of $160\,\mu$m above the nozzle. The amount of nozzle heating increases as the laser approaches the peak gas density at $\sim 160\,\mu$m and then drops steeply as the laser is focused further away from the nozzle. This is caused partially by the reduced solid angle subtended by the nozzle and also the changing gas density profile. Accounting for ion collisions in the gas before they reach the nozzle means that less overall energy is deposited in the nozzle, but the energy per unit mass of heated material increases due to a higher ion stopping power at lower energies.
The estimates presented in this section suggest that nozzle damage is caused primarily by ion energy deposition rather than a resistive current, so nozzle material has little impact on nozzle survival. On the other hand, just as for EMP emitted from solid targets, the discharge current and EMP amplitude and spectrum change when one moves from conducting to dielectric nozzles.
VI Experimental Results
VI.1 Ion Acceleration Experiment on VEGA-3
An experiment was conducted on the VEGA-3 laser system with the aim of using PW laser pulses to trigger laser channelling in inert gases (He, N2, He-N2) and thereby learn more about ion acceleration mechanisms [24, 25, 25].
[…]
Figure 4: Images corresponding to optical probe arrival $\sim 150$ ps after the drive laser in a He gas with a long-focus shock nozzle (a) Raw interferogram (b) Density map showing a plasma channel straddling the peak density region at $x=250$ $\mu$m. The laser focus position in the vacuum was set at $x=0$ $\mu$m.
To measure the electron density in the plasma channel, a folding-wave interferometer was used to take optical interferograms of the ionised gas at different temporal delays relative to the drive laser [57]. Figure 4b shows a He density distribution from the experiment taken $\simeq 150$ ps after the arrival of the laser, revealing a laser channel with average radius $\simeq 70\leavevmode\nobreak\ \mu$m. Modelling the expansion as a Sedov-Taylor cylindrical blast wave [58, 59], $R(t)=\xi_{0}(\mathcal{E}_{b}t^{2}/\rho_{\rm gas}l_{p})^{1/4}$ gives an energy of $\mathcal{E}_{b}\simeq 1.3$ J (taking $\xi_{0}=0.9$ [59]) deposited in the channel over its length of $l_{p}\simeq 0.5$ mm, which is less than 20% the deposited laser energy of 7.3 J. Such a significant difference is explained by the fact that the expanding plasma is essentially collisionless. Plasma ions propagate through the ambient gas and ionise atoms without creating a density compression - only a small portion of the slow ion population contribute to shock wave formation.
2D PIC simulations in CALDER[25, 29] and experimental interferograms of the gas at the point of laser arrival suggest that, by the time the laser has travelled through the low-density wings of the gas density profile and arrived in the central high-density region, the laser has become significantly defocused. For a gas density of 4.5 bar and a plasma channel radius of $20\leavevmode\nobreak\ \mu$m, the plasma volume is $\sim 6\times 10^{-7}$ cm3, the laser energy deposited in the gas is 7.3 J, the bulk electron temperature is $T_{e}\simeq 140$ keV and the pressure is $\sim$50 Mbar (see Tab. 1).
The stopping power of 140 keV electrons in He is $\sim 3$ MeV cm2/g [42], which corresponds to a mean free path $>50$ cm for our gas density. The Debye radius, however, is of order $0.1\leavevmode\nobreak\ \mu$m, so bulk electrons remain in the plasma due to the Coulomb attraction and the plasma expansion is quasi-neutral. Bremsstrahlung losses are negligible on this timescale, so the plasma expands adiabatically.
Assuming that all bulk electron energy is transferred to ions, the average energy of helium ions is about 0.43 MeV, their velocity is $\sim 5\,\mu$m/ps and the expansion time is 100 ps. In this experiment, about 10% of the absorbed laser energy is transferred to hot electrons with an effective temperature of $\sim$1.1 MeV given by the ponderomotive scaling. Then using the plasma charging model described in Sec. II, we estimate a plasma capacitance of 0.0086 pF, a plasma potential of 4.4 MV and 42 nC charge of escaped electrons. The plasma electrostatic energy is about 0.1 J and the escaped electrons have carried away $\sim$0.15 J.
Applying the analysis of gas ionisation described in Sec. II, we find that the photoionisation probability is very low - about $1.6\times 10^{-7}$ the level of ionisation by fast electrons is two orders of magnitude larger at $\sim 3.3\times 10^{-5}$ and the dominant ionisation processes are electrical breakdown and ions in the expanding plasma. The plasma ionisation time is therefore within a few tens of ps. The plasma discharge results in EMP emission with energy $\sim$0.17 mJ, which corresponds to an electric field of $\sim 6.3$ kV/m and a magnetic field of $21\,\mu$T at 1 m from the source. Nozzle damage is due to energy deposition of a few joules by the plasma ions.
Figure 5: Interferograms of the VEGA-3 laser interacting with an N2 gas ejected from a short-focus shock nozzle at different probe times relative to the arrival of the pump beam: (a) Probe time $t\sim 40$ ps (b) Probe time $t\sim 90$ ps.
Progress of gas ionisation between the laser and nozzle was measured with a $2\omega$ pick-off beam directed transversely to the laser channel for interferometric imaging. Figure 5 shows interferograms taken at two different times after the arrival of the laser pulse. The gas is N2, here. Figure 5(a) corresponds to the earliest probing time measured after the laser pulse arrival at $t=0$ ps. The shadow of the gas is visible in both images, suggesting that the entire gas volume can be ionised within 40 ps. The speed of the ionisation front therefore exceeds $\sim 20\,\mu$m/ps, which is consistent with gas ionisation by electrical breakdown.
To monitor ion emission during the experiment, five diamond time-of-flight detectors [60] were placed in multiple different locations around the vacuum chamber ($0^{\circ}$, ±$\pm 17^{\circ}$, $4.85^{\circ}$, $5.5^{\circ}$ with respect to the laser axis, always inclined at $9^{\circ}$ to the horizontal plane of the laser) and a Thomson parabola spectrometer was placed along the laser axis in the same horizontal plane as the laser. When the diamond detectors were placed at ±$\pm 17^{\circ}$, nitrogen ions with energy up to $\sim 80$ MeV were measured. For shots on He gas, $\alpha$-particles were accelerated to 0.7 MeV/u with $10^{9}$ particles per steradian, while shots on N2 gas produced ions up to 5.7 MeV/u with $10^{7}$ particles per steradian. These ions were measured in the laser forward direction (within a ±$\pm 17^{\circ}$ cone)[25].
VI.2 EMP Experiment on VEGA-3
A second experiment was conducted on VEGA-3 with a view to characterising the EMP and ions produced in laser-gas interactions. The laser energy before compressor was $\sim 30$ J and the pulse duration was 30 fs. Based on images of the focal spot taken at low energy, 21% of the energy on-target was contained within the first Airy disk[23] for an on-target intensity of order $\sim 10^{20}$ W cm-2 at best compression, similar to the intensity reported for the experiment in the previous section. The laser contrast was $10^{-12}$ at $0.1$ ns, with no significant pre-pulses[23]. A Prodyn RB-230(R) radiation-hardened B-dot probe was used to measure the amplitude of the magnetic field at different distances from the gas jet, while two diamond time of flight detectors and a Thomson parabola were used to monitor charged particle emission. Signal degradation due to EMP pick-up in the unshielded oscilloscope prevented the easy identification of a photo-peak in the time of flight data, so quantitative analysis of the ion spectrum is not possible. B-dot signals were recorded on a Rohde&Schwarz RTO64 digital oscilloscope with 2 GHz bandwidth and 10 GS/s sample rate. The B-dot probe was oriented so that its axis was parallel to the ground and sensitive to an azimuthal magnetic field relative to the nozzle axis.
B-dot voltage signals were bandpass-filtered between 0.4 GHz and 2 GHz, with the lower limit determined by the probe frequency response [61, 62] and the upper limit by the oscilloscope bandwidth. The waveforms were then cropped, the zero-point offset removed, and cable attenuation corrected in the Fourier domain. The cable attenuation functions were measured using a Rohde&Schwarz ZNA4 Vector Network analyser from 5 MHz to 4 GHz, but the correction was only applied between 0.4 GHz and 2 GHz to avoid amplifying noise beyond the sensitivity range of the diagnostic. Finally, the voltages were integrated in time and multiplied by the probe effective area of 0.2 cm2 (supplied by the manufacturer) to yield the magnetic field.
Figure 6: 3D graphic of the gas jet nozzle and solenoid valve assembly at VEGA-3. Arrows indicate dimensions relevant to electromagnetic emission.
Both solid and gaseous targets were studied during the experiment. The gas was a 3:97 H2-He mix at $\sim 1000$ bar backing pressure, forced vertically downwards through the nozzle aperture to the laser focus. On shots with solid targets, the nozzle and solenoid valve assembly (Fig. 6) was replaced by a bracket and multi-foil array. The foils were made from $6\leavevmode\nobreak\ \mu$m-thick aluminium. Figure 7 shows typical EMP waveforms for solid and gaseous targets, demonstrating a reduction by a factor $2-3$ in the peak magnetic field for the same laser parameters when solid targets were switched to gas. This is consistent with the results of Kugland et al. [8] on the PHELIX laser with an Ar gas jet and tallies with our theoretical estimates. The ChoCoLaT2 code [63] evaluates the charge produced by the VEGA-3 laser interacting with $6\leavevmode\nobreak\ \mu$m-thick Al foils to be of order 100 nC, which is twice higher than the 42 nC estimated for a gas in Sec. VI.1.
Figure 7: Comparison of EMP waveforms for solid and gaseous targets on VEGA-3. The signals were measured using a Prodyn RB-230(R) probe positioned at $60^{\circ}$ to the laser forward direction at a horizontal distance of $2.66$ m from the nozzle and vertically in-line with the laser focal spot. The maximum amplitude of the magnetic field was a factor $2-3$ times lower for the gas targets compared to $6\leavevmode\nobreak\ \mu$m-thick solid Al foils.
Figure 8: Variation of EMP maximum magnetic field with distance from the gas jet. Data was collected with the B-dot probe positioned at $\phi=60^{\circ}$ to the laser axis, with line of sight to the target occluded. The fitted curve is for a $3$ cm-tall antenna with the angle between the antenna axis and the probe assumed constant at $\theta=90^{\circ}$ for the different distances.
VI.2.1 Comparison with EMP Emission Model
Figure 8 shows measurements of the decay of the magnetic field with distance from the gas jet. The data points come from experimental measurements and represent the average of the maximum and minimum magnetic field values in the waveform. Error bars represent one standard deviation from the mean. The green points represent data taken when the B-dot was positioned at $60^{\circ}$ to the laser axis, outside the vacuum chamber and behind a glass window. A Thomson parabola spectrometer inside the chamber blocked the probe’s direct line of sight to the target.
The magnitude of the magnetic field is of order $5-10\,\mu$T at a distance of $\sim 2$ m from the jet. Significant shot-to-shot variations are due to adjustments in the laser focal position inside the gas to optimise the laser coupling to the target. There are insufficient data points to identify an unambiguous scaling with distance, but they can allow us to estimate the charge accumulated in the gas by the laser ejection of hot electrons if we assume a dipolar radiation field. The solid line represents a least squares fit to the data points with Eq. (1) from Ref. [51], assuming an antenna height of $h_{d}=3$ cm (consistent with the vertical height of the jet nozzle in Fig. 6) and leaving the target charge $Q$ as a free parameter. The best-fit charge is $\sim 20$ nC in this case. This is in agreement with the model presented in Sec. II, which predicts a charge of $\sim 40$ nC for the VEGA-3 experimental conditions.
Peaks in the EMP Fourier spectrum constrain the EMP emission mechanism. The average of the EMP Fourier spectra for shots with the gas jet targets reveals multiple prominent resonances between $\nu_{\rm emp}\sim 1.2$ GHz and 1.9 GHz (not shown here). Assuming antenna emission from a monopole of height $c/4\nu_{\rm emp}$, the strongest resonances correspond to heights of $\sim 4-5$ cm - similar to the dimensions of the nozzle. Note that our measurements were only sensitive to emissions from 0.4 GHz to 2 GHz and there were other large metallic objects in the chamber that might contribute to the EMP.
VII Conclusions
We have explored various mechanisms of gas ionisation and nozzle damage in laser-gas interactions. Depending on the gas density and plasma temperature, an electrical discharge and EMP will be triggered either by plasma expansion from the laser focus or electrical breakdown of the ambient gas. The relatively small plasma volume produced in laser-gas interactions leads to a small plasma capacitance and this in turn produces a smaller ejected charge and lower level of EMP compared to solid targets. Nozzle damage is caused by plasma ions depositing energy in the nozzle surface rather than Ohmic heating from the discharge current. Two experiments on the VEGA-3 laser have provided useful supporting evidence for our model. The minimum speed of gas ionisation - inferred from optical interferometry - is consistent with collisional ionisation by ions in an expanding plasma or with the electric discharge.
Further dedicated experiments are needed to confirm the model presented in this paper, in particular by measuring the emitted ion and electron spectra over a large solid angle. Measuring the average ion energy will allow us to estimate the discharge time and plasma expansion velocity. It is also important to measure the laser-ejected charge since 10 times more hot electrons imply 10 times stronger EMP fields. Direct measurement of a discharge current in the nozzle and characterisation of the EMP spectrum is essential to confirm that the EMP is related to antenna emission from the nozzle.
EMP sources couple with the chamber and objects within it, exciting resonant modes[6]. These modes are excited differently according to the nature of the driver and may last much longer than the source itself, depending on their relaxation time. We have discussed a plasma discharge as one important source of EMP, but the electron beam ejected from the plasma can also induce currents and secondary electromagnetic fields in the nozzle and surrounding chamber [64]. The intensity of the induced current is much smaller than that of the escaped electrons, however, and its impact on the nozzle damage is expected to be small.
Experimental validation of the nozzle damage mechanism requires measurement of the ion energy distribution in the radial direction, in the energy range below 1 MeV. This might also be combined with a study of nozzle damage as the laser is focused at various different heights above the gas nozzle. Switching from metallic to ceramic nozzles would be useful for assessing the impact of nozzle material.
Acknowledgements.
The authors would like to thank the staff at the VEGA-3 Laser Facility for their hard work over the course of several experiments. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No. 101052200 — EUROfusion). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. We acknowledge GENCI for providing us access to the Joliot-Curie supercomputer (grants 2021-A0130512993 and 2022-A0130512993). We also acknowledge Grant PID2021-125389OA-I00, funded by MCIN/ AEI / 10.13039/501100011033 / FEDER, UE and “ERDF A way of making Europe”, funded by the European Union. V. Ospina-Bohórquez and C. Vlachos acknowledge support from the LIGHT S&T Graduate Program (PIA3 Investment for the Future Program, ANR-17-EURE-0027). This work received funding from the European Union’s Horizon 2020 research and innovation program through the European IMPULSE project under grant agreement No 871161 and from LASERLAB-EUROPE V under grant agreement No 871124. It is published as part of the international project “PMW”, co-financed by the Polish Ministry of Science and Higher Education within the framework of the scientific financial resources for 2021-2022 under the contract no 5205/CELIA/2021/0 (project CNRS no 239915). Finally, we acknowledge the financial support of the IdEx University of Bordeaux / Grand Research Program “GPR LIGHT”.